OpenStudy (anonymous):

HELP FOR MEDAL!! @misty1212 In Honolulu, Hawaii, the forecast for the next three days is for high temperatures of 80°, 81°, and 80°. 1st objective: What would the temperature on the fourth day need to be to have a 4-day average of °87? 2nd objective: The perimeter of the Hawaiian island of Oahu is approximately 230 miles. If the island were a rectangle with a length that is 35 miles longer than its width, write a system of two-variable equations that you could use to find its length. 3rd objective: Solve the system of equations using substitution or linear combination. Show

OpenStudy (misty1212):

HI again!!

OpenStudy (anonymous):

hi thanks for coming to help me!

OpenStudy (misty1212):

no problem!

OpenStudy (misty1212):

i guess it would have to be pretty hot to get a four day average of 87 we can solve it by calling it \(x\), taking the average\[\frac{80+81+80+x}{4}\] and setting it equal to \(87\) to see what we get

OpenStudy (misty1212):

in other words, solve \[\frac{80+81+80+x}{4}=87\] first add all that stuff

OpenStudy (misty1212):

i get \[\frac{241+x}{4}=87\] so multiply by \(4\) and get \[284+x=348\]

OpenStudy (misty1212):

subtract and you are done

OpenStudy (anonymous):

is x 64 @misty1212?

OpenStudy (misty1212):

i doubt it

OpenStudy (misty1212):

what is \(348-241\)?

OpenStudy (anonymous):

107?

OpenStudy (misty1212):

oh i see i made a typo sorry

OpenStudy (misty1212):

yeah seems hot doesn't it? lets check the answer

OpenStudy (misty1212):

compute \[\frac{80+81+80+107}{4}\] and see if you get \(87\)

OpenStudy (misty1212):

i do!

OpenStudy (misty1212):

ready for number two? The perimeter of the Hawaiian island of Oahu is approximately 230 miles. If the island were a rectangle with a length that is 35 miles longer than its width, write a system of two-variable equations that you could use to find its length.

OpenStudy (anonymous):

yep how do I do that @misty1212?

OpenStudy (misty1212):

it wants a system of two equations, so i guess we have to do that

OpenStudy (misty1212):

the perimeter is \(2l+2w\) so one equation is \[2l+2w=230\]

OpenStudy (misty1212):

then since the length is 35 miles longer than the width, you have \[l=w+35\]

OpenStudy (misty1212):

ooops typo the two equations are \[2l+2w=230\\ l=w+35\]

OpenStudy (misty1212):

you good with that?

OpenStudy (anonymous):

yep, I see it makes sense @misty1212

OpenStudy (misty1212):

ok so on to 3rd objective: Solve the system of equations using substitution or linear combination. Show

OpenStudy (misty1212):

since we know \(l=w+35\) we can substitute it in to the top equation and turn ' \[2l+2w=230\] in to \[2(w+35)+2w=230\] and solve for \(w\)

OpenStudy (misty1212):

here are a few ways to solve this i would start by dividing by 2 to get \[w+35+w=115\] or \[2w+35=115\] and solve that one

OpenStudy (anonymous):

hmm w=35 too @misty1212?

OpenStudy (anonymous):

oh no never mind it's 40

OpenStudy (anonymous):

thanks again for the help @misty1212!